Circular No. 8172

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 O2 (LINEAR) M. Bezpalko, Lincoln Laboratory, reports the LINEAR discovery of

a comet (discovery observation below), showing a tail ≈ 42′′ long in p.a. 230^{o}. Other CCD observers report mag 16.9–17.9 and a tail of up to 6′ long in p.a. 245^{o}–250^{o} on July 30–31 (including S. Sanchez, R. Stoss, and J. Nomen at Mallorca; R. Trentman and R. Frederick at Louisburg, KS; and P. Birtwhistle at Great Shefford, U.K., who also noted a 9′′ central condensation of mag 17.9, adding that the tail was very diffuse and wide).

2003 UT  α_{2000}  δ_{2000}  Mag.
July 29.37618  1^{h}55^{m}23.^{s}89  +14^{o}28′46′′.1  18.6

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-O43. The comet is likely to be of short orbital period.

T = 2003 Sept. 2.110 TT  ω = 30.^{o}140
Ω = 347.894 }2000.0
q = 1.60716 AU  i = 15.623

SUPERNOVA 2003gt IN NGC 6930 Further to IAUC 8169, J. Graham and W. Li report the LOTOSS

discovery of an apparent supernova on an unfiltered KAIT image taken on July 29.4 UT (mag ∼ 16.4). The new object was confirmed on July 31.4 under poor conditions. SN 2003gt is located at α = 20^{h}32^{m}59^{s}.12, δ = +9^{o}52′19′′.3 (equinox 2000.0), which is 4′′.9 east and 4′′.7 south of the nucleus of NGC 6930. A KAIT image taken on July 13.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2003gs IN NGC 936 Further to IAUC 8171, T. Matheson, Harvard-Smithsonian Center for

Astrophysics; and N. Suntzeff, CTIO, report that a detailed inspection of the July 30.4 CTIO spectrum of SN 2003gs shows it to be very similar to a spectrum of the subluminous type-Ia SN 1999by at 1 day before maximum (see Garnavich et al. 2003, Ap.J., in press; http://xxx.lanl.gov/abs/ astro-ph/?0105490). Additional spectroscopic and photometric observations are urged. The redshift mentioned on IAUC 8171 is with respect to the host galaxy.

2003 July 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green